AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 17:18:34 2025
Author
*
:
Subject
*
:
<p>Detector is not biasing as it has in previous tests. Last bias 15/1/22 see previous elogs, max leakage current ~10 uA at 150V</p> <p>Detector bias increases well above 10uA for voltages as low as 30-40V. Optical windows covered, Y09 valves closed, large ion pump OFF and both ion gauges OFF produces similar results but with a slightly lower leakage current.</p> <p>Grounding and pulser cables removed from FEEs and jumper attached to nn bias adaptor card - no change. Bias module swapped out - no effect. NIM rack shows expected voltages (6,12,24) on multimeter.</p> <p> </p> <p>AIDA03 - disconnected. Behaves as expected. Get waveforms and pulser peak. Pulser peak FWHM ~ 70 channels, noise transients observed.</p> <p>AIDA02 - Not producing ADC data</p> <p>AIDA01,04,05,06,07,08 - Very high rates - noisy</p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown